OpenStudy (anonymous):

Stefan left school and traveled toward his friend's house. Eugene left two hours later traveling at 40 km/h in an effort to catch up to Stefan. After traveling for three hours Eugene finally caught up. What was Stefan's average speed?

OpenStudy (anonymous):

need help

OpenStudy (anonymous):

Ok, ready to work toghether?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok. First Stefan leaves right? Then 2 hours pass by, then Eugene goes drives for 3 hours. If his velocity is 40 km/h and he drove 4 hours, how many Km did he travel?

OpenStudy (anonymous):

sorry, he drove 3 hours

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

160

OpenStudy (anonymous):

120 for 3 hour

OpenStudy (anonymous):

actually 120 V=d/t d=V*t d= 40*3=120

OpenStudy (anonymous):

ok, so if both of them started at the same place, we know they travelled the same distance

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

so Stefan's average speed is 40km/h?

OpenStudy (anonymous):

So, the total time since stefan started the trip, is the first 2 hours plus the 3 hours eugene drove afterwards for a total of 5 hours

OpenStudy (anonymous):

using the formula V= distance / time V= 120 km / 5 hours V= ?

OpenStudy (anonymous):

24

OpenStudy (anonymous):

24

OpenStudy (anonymous):

there you go

OpenStudy (anonymous):

ok thank you